OpenStudy (august899):

Which of the following functions would not model exponential decay? (Multiple Choice)

OpenStudy (august899):

A. h(x)=44(0.89)^2x B. j(x)=(0.34)^x-3 C. b(x)=876(1)^x D. z(x)=7(0.999)^4x-2

OpenStudy (unklerhaukus):

one of the options plots a function that does not decay